OpenStudy (anonymous):

The width of a rectangle is 2ft less than the length. The area is 3ft^2. find the length and width. if you just show me how to get started please

OpenStudy (diyadiya):

let l be the length width =l-2 a= l * w 3= l(l-2)

OpenStudy (anonymous):

i still cant come up with the formula

OpenStudy (diyadiya):

l^2-2l-3=0

OpenStudy (diyadiya):

(l-3)(l+1)=0

OpenStudy (diyadiya):

should i continue ?

OpenStudy (anonymous):

yea...im just reading to see how to come to a completion for the future

OpenStudy (diyadiya):

l-3 =0 l+1=0 l=3 l=-1 (length cant be negative) so length = 3

OpenStudy (diyadiya):

width= length -2 =3-2 =1

OpenStudy (anonymous):

Solve the following for L and w:\[\{w\text{ = }L-2,L* w\text{ = }3\}\]

OpenStudy (anonymous):

ok so it cant never be negative

OpenStudy (anonymous):

if my area is 3^2 so what my width is 1?

OpenStudy (anonymous):

The solution is\[\{\{L\to -1,w\to -3\},\{L\to 3,w\to 1\}\} \]The positive solution is the one that makes physical sense.

OpenStudy (anonymous):

length =3 width = 1

OpenStudy (anonymous):

Correct. Thank you for the medal.

OpenStudy (anonymous):

you are welcome...i appreciate the help from all..hopefully I will get better so I can help someone